OpenStudy (anonymous):

Prove: If n ≡1 (mod 6), then n^2 + 2^n is composite. @MIT 18.02 Multiva…

OpenStudy (agreene):

Given: n ≡1 (mod 6) Prove: n^2+2^n is a composite number. Since: n is defined, and is 1 (mod 6): [1 (mod 6)]^2+2^[1 (mod 6)]= 3 And since 3 is prime: Lemma 1 fails via counter-proof. QED

OpenStudy (anonymous):

oh, i see, so the proposition was false?

OpenStudy (agreene):

It certainly appears so

OpenStudy (anonymous):

lol! Ok, thanks! :)

OpenStudy (agreene):

1 (mod 6) =1 so, 1^2=1 2^1=2 1+2=3 3 is prime, not composite.

OpenStudy (nikvist):

\[n=6m+1\quad,\quad m=1,2,3,\dots\]\[n\equiv 1(mod \,\,6)\quad\Rightarrow\quad n^2\equiv 1(mod \,\,6)\]\[n\equiv 1(mod \,\,6)\quad\Rightarrow\quad 2^n\equiv 2(mod \,\,6)\quad,\quad n-\mbox{odd number}\]\[\Rightarrow\quad n^2+2^n\equiv 3(mod \,\,6)\quad\Rightarrow\quad n^2+2^n=6l+3=3(2l+1)\]\[\Rightarrow\quad n^2+2^n\equiv 0(mod \,\,3)\]